OpenStudy (anonymous):

What is the degree measure of an arc 4pi ft. long in a circle of radius 10 ft.?

OpenStudy (anonymous):

4pi=(2pi10(x/360) blah blah blah x=72 degrees

OpenStudy (anonymous):

circumference of the circle=2pir=20pi 20pi constitutes 360deg arc=4pi 4pi arc constitutes xdeg xdeg=(360*4pi)/20pi=360/5=72degree